Johnny found that the large ones totaled 30 more than 3 times the number of small ones. The ratio of the number of small ones to the number of large ones was 1 to 6. How many were large and how many were small?
Answer by edjones(8007)

b=large, s=small
b=3s+30
.
s/b=1/6
6bs/b=6b/6
6s=b
.
6s=3s+30
6s-3s=3s-3s+30
3s=30
3s/3=30/3
s=10
b=60
.
Ed